OpenStudy (anonymous):

Esther and Jim plan to retricea pond to their property. The area of the pond is 44 square meters. What is the radius? Remember, a=3.14r^2 *Use 3.14=22/7

OpenStudy (anonymous):

44=(22/7)r^2 solve for r

OpenStudy (anonymous):

so... r^2 = 44/(22/7)

OpenStudy (anonymous):

r^2=308/22

OpenStudy (anonymous):

r= sqrt(308/22)

OpenStudy (anonymous):

Thats the answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You will need to work backwards to solve this question since you already know the answer (area) is 44 square meters. 44 = (22/7) * r^2 If you multiply both sides by 7/22, you will get rid of the fraction over on the right side of the equation. (7/22) * 44 = (7/22) * (22/7) * r^2 14 = r^2 so, r = sqrt(14)

OpenStudy (amistre64):

how do you retricea pond?